PLEASE ANSWER QUESTIONS CAREFULLY & ANSWERS MUST BE CLEAR TO READ! THANKS!

5. Total, average, and marginal product Happy Cupcake, a cupcake shop, has the following marginal produtve(labeled MP) for its hourly production. 20 Average Product 16 - 0 14 O 12T MP QUANTITY OF LABOR (Workers)

Use the MP curve to fill in the Total Product column in the following table. Total Product Average Product (Cupcakes per worker) Workers (Cupcakes per hour) 10 14 12 10 12 12 4 9 In the previous table, fill in the Average Product column. Then, on the previous graph, use the blue points (circle symbol) to plot the average product (AP) curve at 1, 2, 3, 4, and 5 workers. (Hint: Be sure to plot from left to right.) Marginal product (MP) and average product (AP) curves always intersect at the maximum ▼ of the AP ▼ curve. Grade It Now Save & Continue Continue without saving
Answer #1
Workers Marginal product (Cupcakes per worker) Total product = summation of MP Average product =TP/L
0 - 0 -
1 10 10 10
2 14 (10+14)=24 24/2= 12
3 12 (24+12)= 36 36/3= 12
4 8 (36+8)= 44 44/4=11
5 1 (44+1)=45 45/5=9

By plotting Average product (AP) ,we get the following graph:

16 fy AP 2.

Marginal product (MP) and Average product (AP) curves always intersect at the maximum of the AP curve (as we can see in the graph that MP=AP at 12 when L=3).